A study discussed that the percent purity of oxygen from a certain supplier yields a mean of 99.61% and a standard deviation of 0.08%. If the distribution of percent purity is said to be in normal distribution, what purity value exceed exactly 5% of the population?


1
Expert's answer
2021-05-18T15:34:01-0400

We need to calculate the value of "y" when the "P(Y\geq y)=0.05." This is computed as follows


"P(Y\geq y)=0.05, \\ 1-P(Y\leq y) =0.05,\\ P(Y\leq y)=0.95 ,\\P\left(\frac{Y-\mu}{\sigma}\leq \frac{y-\mu}{\sigma}\right)=0.95, \\ P\left(Z\leq\frac{y-\mu}{\sigma}\right)=0.95,\\P(Z\leq 1.645)=0.95."


From the above, we know that


"\frac{y-\mu}{\sigma}=1.645".


Solving for "y" we have


"y=(1.645)(\sigma)+\mu".


Substituting the values of "\mu" and

"\sigma" we have the value of "y" to be


"y=(1.645)(0.08)+99.61 \Longrightarrow y=99.7416."


Thus, the purity value that exceeds exactly 5% of the population is "99.7416\%."